OpenStudy (madmerc):

Derive the equation of the parabola with a focus at (-7, 5) and a directrix of y = -11. Derive the equation of the parabola with a focus at (-5, 5) and a directrix of y = -1

OpenStudy (madmerc):

I got the answer to the first one is \[f(x)=\frac{ 1 }{ 32 }(x+7)^2-3\]

OpenStudy (madmerc):

and I know the second one has \[\frac{ 1 }{ 12 } \] and \[(x + 5)^2 + 2\] or \[(x - 5)^2 + 2\]

Nnesha (nnesha):

post the work of one of the question first or 2nd

OpenStudy (madmerc):

@Nnesha (x - x)^2 + (y - -11)^2 = (x - -7)^2 + (y - 5)^2 y^2 + 22y + 121 = (x + 7)^2 + y^2 -10y + 25 32 y = (x + 7)^2 -96 y = (1/32)(x + 7)^2 - 3

OpenStudy (madmerc):

for the first one ^^

OpenStudy (madmerc):

as for the second one I know that because all the answers have those thing(s) in them

OpenStudy (madmerc):

@Nnesha please help me with the second one

Nnesha (nnesha):

2nd would be the same and the equation is to find distance between focus and point on the parabola and distance between directrix and point \[\rm \sqrt{(x-a)^2+(y-b)^2}=\left| y-c \right|\] square both sides \[\rm (x-a)^2+(y-b)^2=(y-c)^2\] where (a,b) is the focus point and c is diretrix

Nnesha (nnesha):

plug in (-5,5) for (a,b) and -1 for c lets see which one u get

OpenStudy (madmerc):

@Nnesha is it 1/12 (x+5)^2 + 2

Nnesha (nnesha):

why not (x−5)^2+2? it would be easy for me to check if u post the work :=))